OpenStudy (anonymous):

Guys I need some help. I will give medal to the person that helps me with this. I need to solve these but honestly I forgot how to. 1.39=a(0)2+b(0)+c 1.26=a(18)2+b(18)+c 0=a(45)2+b(45)+c

OpenStudy (dumbcow):

so you have to find quadratic equation using 3 given points solve system with substitution 1st equation ---> c = 1.39 2nd equation ---> b = (1.26 - 1.39 - 18^2 a)/ 18 3rd equation ---> plug in values for b and c to solve for "a"

OpenStudy (anonymous):

So b would be -18.13?

OpenStudy (dumbcow):

?? you have to find "a" before you know what "b" is

OpenStudy (anonymous):

Crap I just noticed the a there.

OpenStudy (anonymous):

Can you help I never really got this.

OpenStudy (anonymous):

I got about this far. A(45)2+((1.26-1.39-18^2a)/18)(45)+c

OpenStudy (dumbcow):

i already helped, use 3rd equation , after you substitute for b you will have equation in just terms of "a" , then solve linear equation using algebra.....hint it will involve combining like terms and factoring out the "a"

OpenStudy (dumbcow):

good , c=1.396 now distribute the 45 and put constants on 1 side and "a" terms on other side

OpenStudy (dumbcow):

c=1.39 sorry

OpenStudy (anonymous):

45+((.028-.0308-7.2a)/.4)(1)+.0308 Am I on the right track?

OpenStudy (dumbcow):

hmm did you divide everything by 45?

OpenStudy (anonymous):

Yes... Did I do it wrong?

OpenStudy (dumbcow):

only partly wrong....the stuff inside parenthesis is multiplied by 45 so dividing will only cancel out the 45 leaving the stuff inside parenthesis the same

OpenStudy (dumbcow):

\[0 = 45a + (\frac{1.26-1.39-18^{2}a}{18}) + \frac{1.39}{45}\]

OpenStudy (anonymous):

Oh, ok.

OpenStudy (anonymous):

45+(.07-.077-18a)+.0308?

OpenStudy (dumbcow):

yes, be careful with rounding error with decimals....i like to leave it in fractions until the very end

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

gtg I got it from here, thanks for your help.

OpenStudy (dumbcow):

yw

OpenStudy (dumbcow):

here is solution to check your answers: http://www.wolframalpha.com/input/?i=0+%3D+45%5E2a+%2B45b+%2Bc+%2C+1.26+%3D+18%5E2a+%2B18b+%2Bc%2C+1.39+%3D+c

OpenStudy (anonymous):

Convert the decimal numbers to integer fractions, simplify.\[\{0=2025 a+45 b+c,126=100 (324 a+18 b+c),139=100 c\} \]Solve the above for a, b and c.\[\left\{a=-\frac{71}{81000},b=\frac{77}{9000},c=\frac{139}{100}\right\} \]The above convert to five significant digit decimal numbers.\[{-0.00087654, 0.0085556, 1.3900} \]

OpenStudy (anonymous):

Sorry, convert should have been spelled converted.

OpenStudy (anonymous):

(.40855, -7.3611, 1.39) This is what I got, but I don't think it is right.

OpenStudy (anonymous):

Can someone walk me through this please.

OpenStudy (phi):

robtobey posted the answer Here is one way to finish. so far we know c= 1.39 use that value of c in the other two equations, and write them like this: \[ 18^2 a + 18 b + 1.39= 1.26 \\ 45^2 a +45 b +1.39=0 \]

OpenStudy (phi):

I would add -1.39 to both sides of each equation , so we have only unknowns on the left side First equation \[ 18^2 a + 18 b = 1.26-1.39 \\ 18^2 a + 18 b = -0.13 \] and second equation \[ 45^2 a +45 b =-1.39 \]

OpenStudy (phi):

divide the first equation by 18 (all terms both sides) divide the second equation by 45 (all terms, both sides)

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

18a+b=-.00722 45a+b=-.0308

OpenStudy (phi):

ok, but I hope you keep all the decimals for the right hand side. I save them as a variable in the calculator. We want to be accurate here. next subtract one equation from the other. The b- b will "drop out", leaving an equation with only "a" that we can solve.

OpenStudy (anonymous):

-27a=-0.03811111111111111111111111111111

OpenStudy (phi):

the left side looks ok. right hand side should be -.00722- (-.0308) (you don't have to type all the decimals... just enough so I can check the answer... but you do want to use all of them in the calculator)

OpenStudy (anonymous):

ok, I forgot about that . 0.02366

OpenStudy (phi):

you are doing 18a+b=-.00722 45a+b=-.0308

OpenStudy (phi):

now you have -27 a = 0.02366 divide both sides by -27 to find a

OpenStudy (anonymous):

I got this, is this right? -8.7654

OpenStudy (phi):

the calculator should read something like -8.7654 EXP -04 or show a small -04 to the right. that is scientific notation for \[ -8.7654 \cdot 10^{-4} \] if you write it as a number, move the decimal point 4 digits to the left you would get -0.00087654

OpenStudy (anonymous):

Oh, gotcha. Ok so a=-.00087654

OpenStudy (phi):

now solve for b using one of the two equations. The first one is as good as any 18a+b=-.00722 b= -.00722 - 18a

OpenStudy (anonymous):

b=-0.00855

OpenStudy (anonymous):

Ok, thank you sooooooooo much.

OpenStudy (anonymous):

I would give you a medal but I can't.

OpenStudy (phi):

I got the same number for b , but positive